OpenStudy (anonymous):

How do I write 4b − 21 ≥ 3 as an interval notation?

OpenStudy (anonymous):

First you need to solve for b, so start off by adding 21 to both sides

OpenStudy (jiteshmeghwal9):

Add 21 both sides & then divide 4 both sides :)

OpenStudy (jiteshmeghwal9):

& then tell me wht do u gt as \(b\) ?

OpenStudy (anonymous):

6... so I would write it as 6, ???

OpenStudy (anonymous):

you get \[b \ge 6\] that means 6 and everything greater than it so \[[6, \infty)\]

OpenStudy (anonymous):

You are my hero... I have a few more like this that I am SERIOUSLY struggling with... PLEASE stay here and help me.. please please please... I am begging you

OpenStudy (anonymous):

\[4(z-3) \le 3(z+5)\]